# A Survey on 3D Egocentric Human Pose Estimation

### v1

@@ Abstract
Index Terms- 3D egocentric human pose estimation, deep learning models, datasets, evaluation metrics.
@@ 1 Introduction
Human pose estimation has been a longstanding challenge in computer vision with the goal of accurately determining a person’s body pose [1, 2, 3, 4]. It has gained prominence due to its relevance in numerous applications, ranging from animation and gaming to surveillance, healthcare, and human-computer interaction. In the current era dominated by XR-technology and the evolution of human-computer interaction, there exists an increasing demand for precise methodologies to track and comprehend human movements. This need is particularly enhanced when considering the perspective of camera wearers, whose viewpoints offer unique insights into the dynamics of human motion.
Egocentric pose estimation is a specialized subset of 3D human pose estimation that focuses on estimating the pose of the person from the point of view of a wearable camera or device worn by the person (first person perspective). Unlike traditional pose estimation, which relies on external cameras or sensors, egocentric pose estimation offers a unique and immersive perspective on human body representation. Real-time processing, adaptability to different environments, user interaction mechanisms, including gestures, and semantic scene understanding contribute to the effectiveness of egocentric pose estimation systems. Figure 1 shows a visual demonstration of the difference between traditional 3D pose estimation and egocentric 3D pose estimation.
The rise of wearable technology, including smart glasses, body-mounted cameras, and head-mounted displays has significantly fueled interest in egocentric pose estimation. These devices have given researchers and developers unprecedented access to the wearer’s field of view. Through this technology, users can interact with their digital environment in a natural and intuitive manner, enhancing the overall user experience. Egocentric pose estimation plays a crucial role across various domains, facilitating intuitive interactions and insights. Its applications encompass Human-Computer Interaction (HCI) for gesture recognition, enhancing augmented and virtual reality experiences by tracking body movements, healthcare for precise therapy monitoring, biomechanical analysis in sports training, and enhancing realism in professional simulations through accurate movement replication.
This survey aims to explore the multifaceted aspects of egocentric human pose estimation. We delve into the fundamental principles, state-of-the-art methodologies, and the latest advancements in this emerging field. By doing so, we seek to provide a comprehensive overview for researchers and practitioners, offering valuable insights into the opportunities and challenges presented by egocentric pose estimation.
1.1 Challenges
Egocentric pose estimation comes with several challenges, stemming from the complexity of accurately capturing and interpreting human movements from the first-person perspective. Some of the key challenges include:
@@ 1.1 Challenges
  • Dataset Constraints: In-the-wild datasets are crucial for capturing real-world complexity, encompassing variations in lighting, backgrounds, activities, and environments. However, the scarcity of such datasets hinders model generalization, particularly in dynamic environments where users engage in various activities and encounter unpredictable situations. The limited availability of diverse samples in training datasets, often developed on motion capture systems, poses challenges for models aiming to handle real-world complexities in outdoor applications.
1.2 Scope of the Survey
Currently, there are numerous systematic surveys related to 2D and 3D human pose estimation on traditional and deep learning based approaches [7, 8, 9, 10] as well shape recovery based approaches [11, 12]. While comprehensive reviews on hand pose [13] and action recognition [14] from egocentric vision are present, it is noteworthy that, to the best of our knowledge, no comprehensive survey on full body egocentric 3D pose estimation methods has been published to date. This absence underscores a notable gap in existing research, despite the increasing interest and advancements in this domain. Hence, there is a pressing need for a comprehensive survey to offer a consolidated overview of state-of-the-art methodologies, challenges, and future directions in egocentric pose estimation.
In this paper, we first discuss different scopes and challenges of egocentric 3D human pose estimation. In section 2, different widely used datasets in pose estimation from egocentric visions have been illustrated. As discussed in section 3, egocentric pose estimation is divided into 2 main categories on the basis of output generation: skeletal based methods and human model based methods. Skeletal methods explore different egocentric models which are mostly regression based (estimation of 3D joint co-ordinates) and heatmap based (estimation of 2D heatmaps). On the other hand, Model based methods mainly generate human models using different shape recovery methods. Additionally, we present a comprehensive evaluation of egocentric pose estimation models, showcasing various evaluation metrics in Section 4, a detailed performance analysis of state-of-the-art approaches on prominent datasets in Section 5. Lastly, we discuss the limitations and future research scopes for egocentric 3D pose estimation in Section 6.
@@ 2 Datasets
Large scale dataset is one of the key factors in visualizing and analysing a computer vision problem. Though there are many benchmark suite datasets like MPII Human Pose [15] and Human 3.6M [16] for traditional pose estimation, a notable gap exists in the availability of widely applicable benchmark dataset for egocentric pose estimation. Different proposed egocentric pose estimation methods have come up with their own datasets, most of which are constrained to motion capture systems in lab environments. In this section, we cover some popular datasets for 3D pose estimation in egocentric setup. Figure 2 showcases sample images from four different datasets. Table 1 lists the most popular datasets we found for egocentric pose estimation and also provides a concise overview of key features in each of them. Details for each dataset are provided in the text below.
@@
EgoCap [21] proposed a method for creating large training datasets for body-part detection using a marker-less motion capture system. They leveraged eight fixed cameras to estimate 3D skeleton motion. They projected it onto fisheye images from a head-mounted camera setup, enhancing the dataset with background replacement, clothing color variations, and simulated lighting changes. The dataset is augmented using intrinsic recoloring and symmetry, doubling its size. The training set includes around 75,000 annotated fisheye images from six subjects, with two additional subjects for validation. This innovative approach minimizes manual annotation and improves model generalization.
Mo2Cap2 [22] dataset addressed the challenge of acquiring annotated 3D pose data for training deep neural networks, specifically tailored to an egocentric fisheye camera setup. Manual labeling in 3D space is impractical, so marker-less multi-view motion capture is proposed for obtaining 3D annotations. However, obtaining a large, diverse set of annotated egocentric training examples using professional motion capture systems is time-consuming. To overcome this, the authors introduced a synthetic training corpus by rendering a human body model from an egocentric fisheye view. The dataset is built upon the SURREAL [23] dataset, incorporating a variety of motions and body textures. Realistic training images are generated by mimicking real-world camera, lighting, and background scenarios. The synthetic dataset is comprised of 530,000 images with ground truth annotations of 2D and 3D joint positions, providing a diverse and extensive resource for training deep neural networks in egocentric pose estimation.
xr-EgoPose [24] provides an extensive collection of 383,000 frames featuring individuals showcasing a rich diversity of skin tones, body shapes, clothing styles, set with various backgrounds and lighting scenarios. Scenes are randomly generated from mocap data, featuring realistic body types like skinny short to full tall versions and skin tones from white to black. The dataset offers high resolution 8-bit images at 30 fps, accompanied by metadata on 3D joint positions, character height, and more. Prioritizing photorealism, the synthetic dataset is created through Maya animation with mocap data and V-Ray’s physically based rendering setup.
In EgoGlass [25], the authors introduced a new dataset for egocentric human pose estimation captured using their EgoGlass, a wearable device resembling eyeglasses. The dataset is obtained through a capture system comprising the EgoGlass helmet for egocentric views and six external cameras for outside-in motion capture. The EgoGlass frame incorporates lightweight Raspberry Pi spy cameras to address visibility limitations posed by the user’s nose or cheek. The authors used an outside-in motion capture setup with six cameras and employ the OpenPose [2] method for third-person view human pose estimation. The ground truth 3D pose is then projected onto the coordinate system of each body cam mounted on the eyeglass frame.
Dataset Year No. of Images No. of Subjects / Actions Characteristics
@@
EgoCap [21] 2016 100000 8 subjects outside-in; marker-less motion capture system; annotated.
Mo2Cap2[22] 2019 53000 images 3000 actions annotated; 700 different body textures.
@@
xr-EgoPose [24] 2019 383k frames 23 male and 23 female subjects; 9 actions synthetic; scene is generated from randomized characters, environments, lighting rigs and animation.
EgoGlass [25] 2021 173577 frames 10 subjects; 6 actions outside-in capture; every frame contains two views, each captured by a body cam attached to the eyeglass frame.
@@
EgoBody [26] 2022 219,731 frames 15 indoor scenes; 36 subjects two subjects (camera wearer and interactee) involved in different interaction scenarios.
@@
EgoPW [18] 2022 318k frames 10 subjects; 20 actions in-the-wild real data; 20 different clothing styles.
@@
UnrealEgo [17] 2022 900000 images 17 subjects; 30 actions 450k in-the-wild stereo views; Motions, 3D environments, spawning human characters.
@@
EgoGTA [19] 2023 320k frames 101 different actions synthetic; based on GTA-IM containing different daily motions and scene geometry.
@@
ECHP [27] 2023 30 video sequences; 75000 frames 9 subjects; 10 daily actions indoor and outdoor; real-world data.
First2Third-Pose[28] 2023 1950 video sequences for 2.5 hours at 25 fps. 14 subjects; 40 actions indoors and outdoors environments, activities include sports actions and daily tasks.
@@
Ego-Exo4D [29] 2023 5625 video sequences for 1422 hours. 839 subjects; 43 actions 131 different scenes in 13 different cities; comprises skilled human activities (e.g., sports, music, dance, bike repair).
Table 1: Popular datasets for egocentric 3D pose estimation.
@@
EgoBody [26] presents an unique large-scale egocentric dataset capturing 2-person interactions using a Microsoft HoloLens2 headset. Unlike previous datasets, it provided synchronized multi-modal data, including RGB, depth, head, hand, and eye gaze tracking. With 125 sequences from 36 subjects in 15 scenes, it offered accurate 3D human shape, pose, and motion ground-truth. The dataset aims to explore the relationships between human attention, interactions, and motions, overcoming limitations of prior datasets and advancing sociological and human-computer interaction research.
@@
The EgoPW [18] dataset, a groundbreaking contribution, is the first in-the-wild human performance dataset captured by synchronized egocentric and external cameras. It features 10 actors, 20 clothing styles, and 20 actions from 318,000 frames organized into 97 sequences. The dataset introduced 3D poses as pseudo labels, enhancing its significance for in-the-wild 3D pose estimation research.
@@
The UnrealEgo [17] dataset is introduced for robust egocentric 3D human motion capture (MoCap). The dataset includes 17 diverse 3D human models exhibiting various motions in 14 realistic environments. With more than 45,000 motions, UnrealEgo outperformed existing mocap datasets. Stereo fisheye images and depth maps are rendered with high motion variety, capturing complex activities like breakdance and backflips. The dataset provides metadata, including 3D joint positions and camera information, and comprises 450,000 in-the-wild stereo views. UnrealEgo showcases motion diversity through wider joint position distributions compared to xR-EgoPose [24], making it a valuable resource for comprehensive egocentric 3D human motion research. Furthermore, the authors introduced two more benchmark dataset UnrealEgo2 and UnrealEgo-RW (RealWorld) [30] datasets which offer a much larger number of egocentric stereo views with a wider variety of human motions than the existing datasets.
@@
The EgoGTA [19] dataset is introduced for accurate human pose and scene geometry ground truth in training, leveraging the diverse daily motions and ground truth scene geometry of GTA-IM. The methodology involves fitting the SMPL-X [31] model to 3D joint trajectories from GTA-IM, followed by attaching a virtual fisheye camera to the forehead of the SMPL-X [31] model for generating synthetic images, semantic labels, and depth maps with and without the human body. The dataset comprises 320,000 frames across 101 sequences with distinct human body textures.
@@
ECHP [27] dataset addresses 3D human motion estimation from egocentric videos. The dataset includes 65000 training images with third-person and egocentric views, 10k validation images, and a test set with egocentric images and 3D ground truth from VICON Mocap. Egocentric poses are the primary focus. OpenPose [2] extracts 2D joints, and a pretrained model handles human segmentation. Cameras are well-calibrated and Aruco marker [32] aided in obtaining egocentric camera pose. The dataset comprises 30 sequences with 9 subjects and 20 textures performing 10 actions in diverse indoor/outdoor scenes. The test set features 4 unseen subjects, emphasizing generalization. The VICON system captured 17k frames with 3D ground truth in the same indoor setting.
The First2Third-Pose [28] dataset presents a diverse collection of short videos capturing 14 individuals engaged in 40 activities. Utilizing head-mounted cameras for egocentric views and static cameras for side and front perspectives, the dataset ensures a comprehensive exploration of human poses in indoor and outdoor settings. Synchronization is achieved through participants clapping before each activity, offering aligned views for enhanced analysis. With 1950 activity sequences lasting 8 to 25 seconds, this dataset addresses the need for synchronized views and facilitating a deeper understanding of actions in both first and third-person perspectives.
@@
Ego-Exo4D [29] is a groundbreaking multimodal dataset and benchmark suite, offering the largest public collection of time-synchronized first and third-person videos captured by 839 individuals across 131 scenes in 13 cities. This unique dataset is comprised of 1,422 hours of video, featuring both egocentric and multiple synchronized exocentric views. Notably, it introduces the “EgoPose” benchmark, focusing on recovering 3D body and hand movements from egocentric videos. The task involves estimating 17 3D body joint positions and 21 3D joint positions per hand, following the MS COCO convention. Ego-Exo4D has advanced the research in ego-exo video analysis and 3D pose estimation from egocentric perspectives.
@@
In this section, we discuss and review different egocentric pose estimation techniques by classifying them into two categories, namely skeletal and model based approaches. Skeletal-based 3D pose estimation methods [33, 34, 35] leverage the human skeleton’s representation to accurately track and infer 3D joint position and body movements. Model based human body pose estimation utilizes a parametric model to accurately predict and estimate 3D joint locations and articulations from 2D image data, facilitating versatile and realistic pose estimation across diverse body shapes and poses. There has been a growing interest towards 3D mesh recovery of human bodies [36, 37, 38] especially after the release of SMPL [39] and SMPL-X [31] models. The retrieval of human body meshes is pivotal in supporting subsequent tasks like reconstructing clothed humans [40, 41], rendering [42], and modeling avatars [43, 44]. The sub-sections below expand on the different methods in each category. We have further sub-categorized the methods based on some significant features, as highlighted in bold.

### v2

@@ 1 Introduction
Human pose estimation [62, 10, 55, 19] has gained prominence due to its relevance in numerous applications, ranging from animation and gaming to surveillance, healthcare, and human-computer interaction. The rise of wearable technology, including smart glasses, body-mounted cameras, and head-mounted displays has significantly fueled interest in egocentric pose estimation, where the focus is on estimating the pose of the person from the point of view of a wearable camera or device worn by the person (first person perspective). Egocentric pose estimation plays a crucial role across various domains, such as in human computer interaction for gesture recognition, augmented and virtual reality experiences by tracking body movements, healthcare for precise therapy monitoring, biomechanical analysis in sports training, hand-object interaction for contextual understanding, and enhancing realism in professional simulations through accurate movement replication. Unlike traditional pose estimation, which relies on external cameras or sensors, egocentric pose estimation offers a unique and immersive perspective on human body representation. Real-time processing, adaptability to different environments, user interaction mechanisms, including gestures, and semantic scene understanding contribute to the effectiveness of egocentric pose estimation systems. Figure 1 shows the difference between traditional and egocentric 3D human pose estimation.
Challenges for Egocentric 3D Human Pose Estimation stem from the complexity of accurately capturing and interpreting human movements from the first-person perspective. Some of the key challenges include:
@@
  • Dataset Constraints: In-the-wild datasets are essential for capturing real-world complexity, including variations in lighting, backgrounds, activities, and environments. However, their scarcity hinders model generalization, especially in dynamic environments with unpredictable situations. Limited availability of diverse samples, often from motion capture systems, poses challenges for models aiming at real-world outdoor applications.
Scope of the Survey: Currently, there are numerous systematic surveys related to 2D and 3D human pose estimation on traditional and deep learning based approaches [78, 65, 45, 16] as well shape recovery based approaches [40, 59]. While comprehensive reviews on hand pose [7] and action recognition [47] from egocentric vision are present, it is noteworthy that, to the best of our knowledge, no comprehensive survey on full body egocentric 3D pose estimation methods has been published to date. This absence underscores a notable gap in existing research, despite the increasing interest and advancement in this domain.
In this survey, we aim to explore the multifaceted aspects of 3D egocentric human pose estimation, by first describing the widely used datasets in Section 2. Next, in Section 3, we explore the different egocentric estimation methods by dividing them into two categories on the basis of output generation: skeletal based methods and human body shape based methods. Skeletal methods explore different methods which are mostly regression based (estimation of 3D joint co-ordinates) and heatmap based (estimation of 2D heatmaps). On the other hand, body shape based methods mainly generate human models using different shape recovery methods. Additionally, we present a comprehensive evaluation of egocentric pose estimation models, showcasing various evaluation metrics in Section 4 and a detailed performance analysis of state-of-the-art approaches on prominent datasets in Section 5. Lastly, we conclude the survey in Section 6 with some future research scopes for egocentric 3D human pose estimation.
@@ 2 Datasets
Large scale dataset is one of the key factors in visualizing and analysing a computer vision problem. While benchmark datasets like MPII [6] and Human3.6M [26] exist for traditional human pose estimation, there’s a notable gap for egocentric pose estimation benchmark datasets. Figure 2 showcases sample images from 4 different datasets. Table 1 summarizes the key features of 9 egocentric pose estimation datasets, with more details provided in the text below.
@@
EgoCap [51] proposed a method for creating large training datasets using a marker-less motion capture system. They leveraged eight fixed cameras to estimate 3D skeleton motion. They projected it onto fisheye images from a head-mounted camera setup, enhancing the dataset with background replacement, clothing color variations, and simulated lighting changes. The training set includes 75,000 annotated fisheye images from six subjects and 25,000 images from two additional subjects for validation.
Dataset Year No. of Images No. of Subjects / Actions Characteristics Dataset Website
@@
EgoCap [51] 2016 100,000 8 subjects marker-less motion capture system; annotated. Link
Mo2Cap2[68] 2019 530,000 3000 actions annotated; 700 different body textures. Link
@@
xr-EgoPose [60] 2019 383,000 23 male and 23 female subjects; 9 actions synthetic; scene is generated from randomized characters, environments, lighting rigs and animation. Link
@@
EgoBody [74] 2022 219,731 15 indoor scenes; 36 subjects two subjects (camera wearer and interactee) involved in different interaction scenarios. Link
@@
EgoPW [66] 2022 318,000 10 subjects; 20 actions in-the-wild real data; 20 different clothing styles. Link
@@
UnrealEgo [3] 2022 900,000 17 subjects; 30 actions 450k in-the-wild stereo views; Motions, 3D environments, spawning human characters. Link
@@
EgoGTA [67] 2023 320,000 101 different actions synthetic; based on GTA-IM containing different daily motions and scene geometry. Link
@@
ECHP [39] 2023 30 video sequences; 75,000 frames 9 subjects; 10 daily actions indoor and outdoor; real-world data. Link
@@
Ego-Exo4D [18] 2023 5625 video sequences; 1422 hours 839 subjects; 43 actions 131 different scenes in 13 different cities; comprises skilled human activities (e.g., sports, music, dance, bike repair). Link
Table 1: Popular datasets for egocentric 3D human pose estimation.
The Mo2Cap2 dataset [68] tackles the challenge of obtaining annotated 3D pose data and introduces a marker-less multi-view motion capture. To address the time-consuming nature of obtaining diverse egocentric training examples, the dataset includes a synthetic training corpus generated from egocentric fisheye views. Built upon the SURREAL dataset [63], it offers 530,000 realistic training images with ground truth annotations of 2D and 3D joint positions.
xr-EgoPose [60] provides an extensive collection of 383,000 frames featuring individuals showcasing a rich diversity of skin tones, body shapes, clothing styles, set with various backgrounds and lighting scenarios. Scenes are randomly generated from mocap data, featuring realistic body types like skinny short to full tall versions and skin tones from white to black. Prioritizing photorealism, the synthetic dataset is created through Maya animation with mocap data and V-Ray’s physically based rendering setup.
@@
EgoBody [74] captures 2-person interactions using a Microsoft HoloLens2 headset. It provides synchronized multi-modal data, including RGB, depth, head, hand, and eye gaze tracking. With 125 sequences from 36 subjects in 15 scenes, it offers accurate 3D human shape, pose, and motion ground-truth. The dataset aims to explore the relationships between human attention, interactions, and motions, overcoming limitations of prior datasets, and advancing sociological and human-computer interaction research.
@@
The EgoPW [66] dataset is the first in-the-wild human performance dataset captured by synchronized egocentric and external cameras. It features 10 actors, 20 clothing styles, and 20 actions from 318,000 frames organized into 97 sequences, along with the 3D poses as pseudo labels.
@@
The UnrealEgo dataset [3] introduces robust egocentric 3D human motion capture with 17 diverse 3D models and over 45,000 motions in 14 environments. It has stereo fisheye images and depth maps capturing complex activities like breakdance and backflips. With metadata including 3D joint positions and camera details, it comprises 450,000 in-the-wild stereo views, showcasing wider joint position distributions compared to xR-EgoPose [60]. They followed up with the UnrealEgo2 and UnrealEgo-RW datasets [4], which provide more views with diverse human motions.
@@
The EgoGTA [67] dataset comprises of 320,000 frames across 101 sequences with distinct human body textures, by leveraging the diverse daily motions and ground truth scene geometry of GTA-IM [9]. The methodology involves fitting the SMPL-X [50] model to 3D joint trajectories from GTA-IM [9], followed by attaching a virtual fisheye camera to the forehead for generating synthetic images, semantic labels, and depth maps with and without the human body.
@@
The ECHP [39] dataset consists of 65,000 training images, 10,000 validation images, and a test set with egocentric images and 3D ground truth from VICON Mocap. Egocentric poses are extracted using OpenPose [10] and human segmentation. Calibration and Aruco markers [17] aid in obtaining egocentric camera pose. The dataset has 30 sequences with 9 subjects, 20 textures, and 10 actions in various indoor/outdoor scenes. The test set provides generalization with 4 unseen subjects and 17,000 ground truth frames.
@@
Ego-Exo4D [18] is a groundbreaking multimodal dataset and benchmark suite, offering the largest public collection of time-synchronized first and third-person videos captured by 839 individuals across 131 scenes in 13 cities. It is comprised of 1,422 hours of video, featuring both egocentric and multiple synchronized exocentric views. The EgoPose benchmark focuses on recovering 3D body and hand movements from egocentric videos. The task is to estimate 17 3D body joint positions and 21 3D joint positions per hand, following the MS COCO convention.
Skeletal based Methods Year Highlighted Characteristics Dataset Limitations Code/Project Website Link
Egocap [51] 2016 First marker-less motion capture system; utilized pose estimation framework for fisheye views with a ConvNet based body-part detector. EgoCap No real-time prototype. Project
Jiang et al. [28] 2017 Leveraged dynamic motion signatures and static scene structures to infer the invisible pose efficiently. custom Kinect V2 dataset Ambiguity in egocentric inputs due to unpredictable arm poses.
Mo2Cap2 [68] 2019 Real-time; disentangled 3D pose estimation, addressed 2D joint detection, camera-to-joint distances, and joint position recovery for accurate results and a precise 2D overlay. Mo2Cap2 Scenes with severe occlusions. Project
xr-EgoPose [60] 2019 Encoder-decoder model for VR headset images, addressing resolution differences in upper and lower body poses, with a dual-branch decoder preserving uncertainty information. xR-EgoPose Scenes with extreme occlusions and out-of-field view. Code
You2Me [46] 2020 Inferred robust poses by incorporating static scene features, explicit second-person body interactions and utilizing dyadic interactions and dynamic first-person motion features. You2Me Scenerios where camera wearer is crouched and camera points towards the floor. Code
EgoGlass [77] 2021 Utilized body part information for low-visible joints and tackling self-occlusion by preserving uncertainty information. EgoGlass Lower body estimation produces larger errors.
Zhang et al. [76] 2021 Implemented auto-calibration module with self-correction for fisheye cameras to rectify image distortions, ensuring alignment between 3D predictions and distorted 2D poses. xR-EgoPose Not evaluated in real-world setting.
Wang et al. [64] 2021 Spatio-temporal optimization framework that combines 2D and 3D keypoints, VAE-based motion priors and SLAM-based camera pose estimation for stable global body pose estimation. Mo2Cap2, AMASS Not evaluated in real-world setting. Project
Wang et al. [66] 2022 Implemented weak supervision with spatio-temporal optimization and synthetic data with domain adaptation for better egocentric pose estimation. EgoPW Accuracy of pseudo labels constrained by in-the-wild capture system.
Akada et al. [3] 2022 Enhances 3D pose estimation by integrating a stereo-based 2D joint location estimation module with weight-sharing encoders and a multi-branch autoencoder for uncertainty capture. UnrealEgo Occlusions and complex motions scenerios. Project
Ego+X [38] 2022 Dual-camera framework for 3D global pose estimation and social interaction characterization, leveraging visual SLAM and a Pose Refine Module (PRM) for spatial and temporal accuracy and characterizes social interactions based on global 3D poses. ECHA Camera localization robustness limited; temporal smoothing effectiveness not fully evaluated.
Wang et al. [67] 2023 First egocentric pose estimation framework, integrating depth estimation for occlusion handling in close interactions. EgoGTA, EgoPW-Scene Accuracy is constrained by depth estimation where scene is occluded by body. Project
EgoFish3D [39] 2023 A self-supervised framework for egocentric 3D pose estimation, utilizing real-world data with three key modules: third person view, egocentric, and interactive modules, achieving accurate results without the need for ground truth annotations. ECHP Overlooked the significance of the perspective factor, which can convey valuable information about the 3D effect intensity.
Ego3DPose [33] 2023 A stereo matcher network and perspective embedding heatmap representation, independent learning of stereo correspondences and leveraging 3D perspective information. UnrealEgo Scenes with occlusions, distortions and real-world setting. Code
Ego-STAN [48] 2023 Tackles fisheye distortion and self-occlusions in egocentric human pose estimation through a domain-guided spatio-temporal transformer, using 2D image representations, feature map tokens, and 3D pose estimation for accurate joint localization and uncertainty management. xr-EgoPose Scenes in real-world setting.
Dhamanaskar et al. [13] 2023 Utilized third-person view information, creating a self-supervised neural network that establishes a shared space for consistent 3D body pose detection across diverse video settings, ensuring adaptability to real-world scenarios with unknown camera configurations. First2Third-Pose Evaluation limited to two datasets; broader assessment needed for generalization. Code
EgoFormer [36] 2023 Leveraged video context and establishing long-term temporal relationships. It addresses ambiguity in first-person videos, surpassing dynamic features, and introduces a novel motion clue representation for enhanced accuracy. CMU Mocap [1] Lack of real-world testing and limited model comparisons.
@@
After performing an extensive literature search for egocentric pose estimation, in this section, we discuss around 35 popular techniques by classifying them into two categories, namely skeletal and body shape based approaches. Skeletal-based 3D pose estimation methods [49, 56, 57] leverage the human skeleton representation to accurately track and infer 3D joint position and body movements. Human body shape based human body pose estimation methods [8, 73, 23] utilize a parametric model, such as SMPL [41] and SMPL-X [50], to accurately estimate 3D joint locations and body shapes. The retrieval of human body meshes is pivotal in supporting subsequent tasks like reconstructing clothed humans [70, 79], rendering [22], and modeling avatars [24, 80]. The sub-sections below expand on the different methods in each category. We have further sub-categorized the methods based on some significant features, as highlighted in bold.